Julius can paint a house three times faster than Ruben. Working together, they can paint a house in 4 days. How long would it take the faster painter if he works alone?
Answer by mananth(16946)

Julius can paint a house three times faster than Ruben. Working together, they can paint a house in 4 days. How long would it take the faster painter if he works alone?

let Julius take x hours to paint
he does 1/x of job in 1 hour

Rueben takes 3x hours
Rueben does 1/3x of the job in 1 hour
together they take 4 days
they do1/4 of the job in 1 day
1/x + 1/3x = 1/4
4/3x= 1/4
3x=16
x= 16/3 hours
julien Working together, they can paint a house in 4 days. How long would it take the faster painter if he works alone?